OpenStudy (ktbug3):

62%of 89.3 will fan n medal

OpenStudy (mathmale):

what did you learn from the previous probelm we worked on? I'd hope you'd make some effort to solve each problem BEFORE asking for help with it. What is the general formula for finding p percent of some number n? think back: how have you approached similar problems in the past?

OpenStudy (ktbug3):

62/89.3x100% somewhat

OpenStudy (c1ndrella78):

1. We assume, that the number 62 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 100% equals 62, so we can write it down as 100%=62. 4. We know, that x% equals 89.3 of the output value, so we can write it down as x%=89.3. 5. Now we have two simple equations: 1) 100%=62 2) x%=89.3 where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 100%/x%=62/89.3 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. Here's the equation 100%/x%=62/89.3 Do you think you can solve this by yourself?

OpenStudy (gagecool):

The answer is 55.366

OpenStudy (ktbug3):

is it 69.4?

OpenStudy (c1ndrella78):

No, it's not 69.4, try to solve the equation

OpenStudy (ktbug3):

ok

OpenStudy (mathmale):

Wait, please. Do not hand out answers without involvement on the part of the person who asked the question. Cinderella has discussed the procedure for finding a percentage of a given quantity, and has set the groundwork for ktbug3 to find her own answer. THAT is what we're looking for here on OpenStudy. ktbug3? Your move.

OpenStudy (ktbug3):

55.366?

OpenStudy (mathmale):

@Gagecool The answer is 5xxxxx Do NOT hand out answers! @ktbug3: Please show all steps you used to find that result.

OpenStudy (c1ndrella78):

Can you show me how you got that?

OpenStudy (ktbug3):

yes

OpenStudy (mathmale):

@ktbug3: show and explain how you got your answer, please.

OpenStudy (ktbug3):

i cross multiplied 62/100x x/89.3 then got 100x=5536.6 and divided by 100 on bothsides to get 55.366

OpenStudy (ktbug3):

am i correct?

OpenStudy (mathmale):

That's great. Exactly what I was hoping for. Because you have explained what you have done, you will understand this process better next time through. Thank you. :)

OpenStudy (ktbug3):

the x after 100 is ment to be a *

OpenStudy (gagecool):

yes you are right

OpenStudy (ktbug3):

ok great thnx

OpenStudy (mathmale):

Cinderella: Thank you for explaining the process here and for not giving away the answer. Well done.

OpenStudy (gagecool):

okay next time i will explain and give the steps and not the answer i just knew the answer so i just put it down. Wont happen again but your welcome ktbug3

OpenStudy (mathmale):

Thank you, gagecool.